OpenStudy (anonymous):

What is the solution for 4x = 2x + 4?

OpenStudy (ricardovg):

4x = 2x + 4 2x = 4 x = 2

hero (hero):

4x = 2x + 4 4x - 2x = 2x -2x +4 2x = 4 x =4/2 x = 2

OpenStudy (anonymous):

take 2x from both sides of the equation then divide by 2. x=2